OpenStudy (anonymous):

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $123.00 for 3 days and 300 miles, while Mary was charged 226.00 for 5 days and 600 miles. What does Best Rentals charge per day and per mile?

OpenStudy (jackellyn):

Let d=charge per day in $ and m=charge per mile in $ Barney: 3d + 300m = 135 Mary: 5d + 600m = 250 You want to solve for d and m you can do it by substitution or elimination, can you do that?

OpenStudy (anonymous):

yeah i got 4.75 for the charge per mile and 430 for the charge per day which i think is wrong

OpenStudy (jackellyn):

i prefer substitution so lets solve for m in the first eqn 3d + 300m = 135 300m = 135-3d m= (135-3d)/300 substitute that into the second eqn 5d + 600(135-3d/300) = 250 5d + 270 - 6d = 250 -d + 270 = 250 -d = 250 -270 -d = -20 d=20 Substitute into first eqn 3(20)+300m=135 60 + 300m = 135 300m= 75 m=1/4=0.25 So Best Rentals charges $20 per day and $0.25 per mile.

OpenStudy (jackellyn):

Do you understand what I did? Any parts confuse you?

OpenStudy (anonymous):

i usually use eliminated becuase i understand it better but i get what you did